OpenStudy (anonymous):

f(x) = xe^3/x a) Use l'Hospital's Rule to determine the limit as x 0+. b)Use calculus to find the local minimum value. c)Find the interval where the function is concave up. d) Find the interval where the function is concave down.

OpenStudy (anonymous):

Yay right up my alley, love calculus!

OpenStudy (anonymous):

wait is it f(x) = x * e^(3/x)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so rewrite that as: f(x) = e^(3/x)/ (1/x)

OpenStudy (anonymous):

according to l'hopitals the limit of that as it goes to 0, is: (e^(3x^-1))'/(x^-1)'

OpenStudy (anonymous):

hold on.. actually can we go on twiddla?

OpenStudy (anonymous):

http://www.twiddla.com/578019 Go there it's easier for me to write.

OpenStudy (bahrom7893):

As I showed u on twiddla the limit diverges.

OpenStudy (bahrom7893):

Here's the picture..

OpenStudy (anonymous):

what is diverges

OpenStudy (bahrom7893):

Does not exist = goes to infinity, answer is inifinity techincally..

OpenStudy (bahrom7893):

f(x) = xe^3/x 2) Use calc to find the local min: that simply means taking the derivative: f'(x) = x*e^(3/x) oh god I hate it when there's a variable in the power.. go back to that twiddla.. this will be messy.

OpenStudy (anonymous):

that wasnt right for the limit

OpenStudy (bahrom7893):

From the plus side? yes it was, and wolframalpha confirms it. Otherwise your question is wrong.

OpenStudy (anonymous):

no it wasnt right

OpenStudy (anonymous):

im putting it into my online homework it said its wrong

OpenStudy (bahrom7893):

can't be true, check the question then. Look wolframalpha is always true.

OpenStudy (bahrom7893):

okay can i see a screen of ur online hw?

OpenStudy (bahrom7893):

Try undefined, undef, DNE, infinity, infinity symbol, etc..

OpenStudy (anonymous):

f(x) = x*e^(3/x)

OpenStudy (anonymous):

infinity was right

OpenStudy (anonymous):

how do i do the others?

OpenStudy (bahrom7893):

okay let's go back to twiddla..

OpenStudy (anonymous):

u are confusing me!

OpenStudy (bahrom7893):

try x = 3 or x = 0, im too bored to check which one is max and which one is min.

OpenStudy (bahrom7893):

did we get it?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

what is that showing

OpenStudy (anonymous):

The actual limits. It doesn't show you how to work them out.

OpenStudy (anonymous):

yes but it doent show the minimum or concave up or down